### v1

@@ Appendix C Proof of Theorem 1
Therefore, whenever $10\leq c<n-$ 1 $\min_{\sigma^{2}\geq\sigma^{2}_{\epsilon,\delta}}\min_{r}\min_{\bm{\alpha}}\mathrm{MSE}(n,\sigma^{2},r,\bm{\alpha})$ is achieved when $\sigma^{2}\to\infty$. For $c=n-1$ which corresponds to the case of LDP, any $\sigma^{2}\geq\sigma^{2}_{\epsilon,\delta}$ (with the corresponding $r_{*}$) gives the same MSE. Hence, the solution $\sigma^{2}=\sigma^{2}_{\epsilon,\delta}$ and $r_{*}=0$ suffices to reach the MMSE when $c=n-1$.
@@
For any $10\leq c<n-$ 1, $\lim_{\sigma^{2}\to\infty}\min_{r}\min_{\bm{\alpha}}\mathrm{MSE}(n,\mathcal{C},\sigma^{2},r,\bm{\alpha})=O\left(\frac{\sigma^{2}_{\epsilon,\delta}}{(n-c)^{2}}\right)$.
Proof:  Here we prove the claim for the most general case of $c>1$. The proof of other two cases $c=0$ and $c=1$ follow similar steps. Recall that for the case where $|\mathcal{F}|=n$,
$\min_{r}\min_{\bm{\alpha}}\mathrm{MSE}(n,\mathcal{C},\sigma^{2},r,\bm{\alpha})=\frac{1}{1+\frac{(n-c)/d}{\sigma^{2}+r_{*}(n-1)}}$ (127)
@@
$r_{*}=\frac{-(n-2)(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})-\sigma^{2}c}{2(n-1)(c-1)}+\frac{\sqrt{((n-2)(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})-\sigma^{2}c)^{2}+4(n-c-1)\sigma^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})}}{2(n-1)(c-1)}$ (128)
@@
×$\quad\times\left(1+\frac{1}{2}\frac{\frac{c^{2}\sigma^{2}\sigma^{2}_{\epsilon,\delta}}{\sigma^{2}-\sigma^{2}_{\epsilon,\delta}}-\sigma^{2}_{\epsilon,\delta}(n-2)^{2}}{\sigma^{2}(n-c)^{2}}-\frac{1}{4}\frac{1}{2!}\left(\frac{\frac{c^{2}\sigma^{2}\sigma^{2}_{\epsilon,\delta}}{\sigma^{2}-\sigma^{2}_{\epsilon,\delta}}-\sigma^{2}_{\epsilon,\delta}(n-2)^{2}}{\sigma^{2}(n-c)^{2}}\right)^{2}+\dotsc\right)$ (130)
@@
$\qquad\qquad\qquad\left.+\sigma^{2}(n-c)\left(1-\frac{1}{2}\frac{\sigma^{2}_{\epsilon,\delta}}{\sigma^{2}}+\frac{1}{2}\frac{\frac{c^{2}\sigma^{2}\sigma^{2}_{\epsilon,\delta}}{\sigma^{2}-\sigma^{2}_{\epsilon,\delta}}-\sigma^{2}_{\epsilon,\delta}(n-2)^{2}}{\sigma^{2}(n-c)^{2}}+O(\frac{1}{\sigma^{4}})\right)\right)$ (131)
$=\frac{1}{2(c-1)}\left(\sigma^{2}_{\epsilon,\delta}(n-2)-\frac{1}{2}\sigma^{2}_{\epsilon,\delta}(n-c)+\frac{c^{2}\sigma^{2}\sigma^{2}_{\epsilon,\delta}}{2(n-c)(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})}-\frac{\sigma^{2}_{\epsilon,\delta}(n-2)^{2}}{2(n-c)}+O(\frac{1}{\sigma^{2}})\right)$ (132)
@@
$\lim_{\sigma^{2}\to\infty}\min_{r}\min_{\bm{\alpha}}\mathrm{MSE}(n,\mathcal{C},\sigma^{2},r,\bm{\alpha})$
$=\lim_{\sigma^{2}\to\infty}\left(1+\frac{(n-c)/d}{\sigma^{2}+r_{*}(n-1)}\right)^{-1}$ (133)
$=\lim_{\sigma^{2}\to\infty}\left(1\!+\!\frac{(n-c)/d}{\frac{1}{2(c-1)}\left(\sigma^{2}_{\epsilon,\delta}(n-2)\!-\!\frac{1}{2}\sigma^{2}_{\epsilon,\delta}(n-c)+\frac{c^{2}\sigma^{2}\sigma^{2}_{\epsilon,\delta}}{2(n-c)(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})}-\frac{\sigma^{2}_{\epsilon,\delta}(n-2)^{2}}{2(n-c)}+O(\frac{1}{\sigma^{2}})\right)}\right)^{-1}$ (134)
$=\lim_{\sigma^{2}\to\infty}\left(1+\frac{(n-c)^{2}}{d\sigma^{2}_{\epsilon,\delta}}\right)^{-1}$ (135)
=$=\begin{cases}O\left(\frac{d\sigma^{2}_{\epsilon,\delta}}{(n-c)^{2}}\right),&(n-c)^{2}>>d\\ O(1),&otherwise.\end{cases}$ (136)
@@
Next, we derive the optimum noise parameters for the case with dropouts, i.e., any $\mathcal{C}$ and $\mathcal{F}$ with $|\mathcal{F}|<n$. We first note that the DME process is the same as what is considered in the case with no dropouts with a reduced number of users. As the noise parameters of the privacy mechanism ($\sigma^{2}$ and $r$) are defined prior to the aggregation stage at the server, they are optimized for the worst case with the largest number of dropouts, resulting in an increase in the MSE compared to the case with no dropouts. However, the decoder can be optimized for the surviving number of users at the time of aggregation, as aggregation is performed after obtaining all responses from the remaining users. Therefore, the proof of the optimum noise parameters and the decoder consists of two steps, namely, 1) determining the optimum decoder for any given number of surviving users, 2) calculating the optimum noise parameters of the privacy mechanism, considering the worst case dropouts and colluding users.
@@
Step 1 is direct from Proposition 1. Proposition 1 provides the optimum decoder for any fixed $\sigma^{2}$ and $r$ with any set of surviving users $\mathcal{F}\subset[1:n]$, $|\mathcal{F}|\geq t$. Even though all users may not respond when allowing for dropouts, the same conditions in Lemma 1 must be satisfied by the noise parameters to satisfy the strongest form of $(\epsilon,\delta)$-DP, considering the maximum information leakage that occurs when all $n$ users respond. Therefore, the problem to be solved is given by,
$\min_{r,\sigma^{2}}\!\!\max_{\begin{subarray}{c}\mathcal{U}\subset[1:n]\\ |\mathcal{C}|\leq c<t\leq|\mathcal{F}|\leq n\end{subarray}}\!\!\min_{\bm{\alpha}}\mathrm{MSE}(n,\mathcal{F},\mathcal{C},\sigma^{2},r,\bm{\alpha})$ $=\min_{\sigma^{2}\geq\sigma^{2}_{\epsilon,\delta}}\!\!\min_{r}\!\!\!\!\max_{\begin{subarray}{c}\mathcal{U}\subset[1:n]\\ |\mathcal{C}|\leq c<t\leq|\mathcal{F}|\leq n\end{subarray}}\!\!\left(1+\frac{|\mathcal{U}|/d}{\sigma^{2}+r(|\mathcal{F}|-1)}\right)^{-1}$ (137)
$=\min_{\sigma^{2}\geq\sigma^{2}_{\epsilon,\delta}}\min_{r}\left(1+\frac{(t-c)/d}{\sigma^{2}+r(t-1)}\right)^{-1},$ (138)
as $\left(1+\frac{|\mathcal{U}|/d}{\sigma^{2}+r(|\mathcal{F}|-1)}\right)^{-1}$ is decreasing in $|\mathcal{F}|$ for any fixed $\sigma^{2}$ and $r$ that result in a valid covariance matrix, i.e., satisfies $\sigma^{2}+r(n-1)>0$. For any fixed $\sigma^{2}\geq\sigma^{2}_{\epsilon,\delta}$, $\left(1+\frac{(t-c)/d}{\sigma^{2}+r(t-1)}\right)^{-1}$ increases with $r$, and the optimal $r$ is given by the $r_{*}$ in (102). The next step is to solve,
@@
$\sigma^{2}_{opt}$ $=\arg\min_{\sigma^{2}\geq\sigma^{2}_{\epsilon,\delta}}\min_{r}\max_{\begin{subarray}{c}\mathcal{U}\subset[1:n]\\ |\mathcal{C}|\leq c<t\leq|\mathcal{F}|\leq n\end{subarray}}\min_{\bm{\alpha}}\mathrm{MSE}(n,\mathcal{F},\mathcal{C},\sigma^{2},r,\bm{\alpha})$ (139)
@@
$=\arg\min_{\sigma^{2}\geq\sigma^{2}_{\epsilon,\delta}}\left(1+\frac{(t-c)/d}{\sigma^{2}+r_{*}(t-1)}\right)^{-1}.$ (140)
@@
$\sigma^{2}_{opt}=\frac{\sigma^{2}_{\epsilon,\delta}(n^{2}-2n-cn+2)}{(n-c)^{2}}+\frac{\sigma^{2}_{\epsilon,\delta}(n-c-1)(n+c-2nc+t(n+c-2))}{(n-c)^{2}\sqrt{(t-c)(n-t)(n-c-1)}}.$ (141)
@@
$G(\sigma^{2})=\min_{r}\max_{\begin{subarray}{c}\mathcal{U}\subset[1:n]\\ |\mathcal{C}|\leq c<t\leq|\mathcal{F}|\leq n\end{subarray}}\min_{\bm{\alpha}}\mathrm{MSE}(n,\mathcal{F},\mathcal{C},\sigma^{2},r,\bm{\alpha})=\left(1+\frac{(t-c)/d}{\sigma^{2}+r_{*}(t-1)}\right)^{-1}$ (142)
@@
$\sigma^{2}_{cr}$ $=\frac{\sigma^{2}_{\epsilon,\delta}(n^{2}-2n-cn+2)}{(n-c)^{2}}+\frac{\sigma^{2}_{\epsilon,\delta}\lambda\sqrt{(n-c-1)(n-t)(t-c)}}{(n-c)^{2}(n-t)(t-c)}$ (143)
@@
Proof: $G(\sigma^{2})=\left(1+\frac{(t-c)/d}{\sigma^{2}+r_{*}(t-1)}\right)^{-1}$ for any $\sigma^{2}\geq\sigma^{2}_{\epsilon,\delta}$ is a smooth and continuous function in $\sigma^{2}$ as $\sigma^{2}+r_{*}(t-1)$ is smooth and $\sigma^{2}+r_{*}(t-1)>0$ (proof in Section C, Claim 4). Consider the derivative of $G(\sigma^{2})$ with respect to $\sigma^{2}$ denoted by $G^{\prime}(\sigma^{2})$ to find its critical points.
$G^{\prime}(\sigma^{2})=-\left(\frac{-\frac{(t-c)}{d}\left(1+(t-1)\frac{dr_{*}}{d\sigma^{2}}\right)}{(\sigma^{2}+r_{*}(t-1))^{2}}\right)\left(1+\frac{(t-c)/d}{\sigma^{2}+r_{*}(t-1)}\right)^{-2}=0$ (144)
@@
$=K(t,c,\sigma^{2},r_{*})\left(1+\frac{t-1}{2(n-1)(c-1)}\left(-(n+c-2)+\frac{1}{2}\frac{1}{\sqrt{A(y)}}\frac{dA(y)}{dy}\right)\right)$ (147)
where $K(t,c,\sigma^{2},r_{*})=\frac{(t-c)/d}{(\sigma^{2}+r_{*}(t-1))^{2}}\left(1+\frac{(t-c)/d}{\sigma^{2}+r_{*}(t-1)}\right)^{-2}>0$ and
@@
$G^{\prime}(\sigma^{2}_{cr}+\tilde{\delta})$ $=K(t,c,\sigma^{2}_{cr}+\tilde{\delta},r_{*})\left(1+\frac{t-1}{2(n-1)(c-1)}\left(-(n-c-2)+h(\frac{\sigma^{2}_{cr}+\tilde{\delta}}{\sigma^{2}_{\epsilon,\delta}})\right)\right)$ (156)
$=K(t,c,\sigma^{2}_{cr}+\tilde{\delta},r_{*})\left(1+\frac{t-1}{2(n-1)(c-1)}\left(-(n-c-2)+h(y_{cr})+\delta\right)\right)$ (157)
@@
$G^{\prime}(\sigma^{2}_{cr}-\tilde{\delta})$ $=K(t,c,\sigma^{2}_{cr}-\tilde{\delta},r_{*})\left(1+\frac{t-1}{2(n-1)(c-1)}\left(-(n-c-2)+h(\frac{\sigma^{2}_{cr}-\tilde{\delta}}{\sigma^{2}_{\epsilon,\delta}})\right)\right)$ (159)
$=K(t,c,\sigma^{2}_{cr}-\tilde{\delta},r_{*})\left(1+\frac{t-1}{2(n-1)(c-1)}\left(-(n-c-2)+h(y_{cr})-\delta\right)\right)$ (160)
@@
$=K(t,c,\sigma^{2},r_{*})\left(1-\frac{t-1}{n-1}\left(1+\frac{n-2}{(y(n-1)-(n-2))^{2}}\right)\right)$ (164)

### v2

@@ Appendix C Proof of Theorem 1
Therefore, whenever $10\leq c<n-$ 1 $\min_{\sigma^{2}\geq\sigma^{2}_{\epsilon,\delta}}\min_{r}\min_{\boldsymbol{\alpha}_{\mathcal{U}_{all}}}\mathrm{MSE}(n,\mathcal{U}_{all},\sigma^{2},r,\boldsymbol{\alpha}_{\mathcal{U}_{all}})$ is achieved when $\sigma^{2}\to\infty$. For $c=n-1$ which corresponds to the case of LDP, any $\sigma^{2}\geq\sigma^{2}_{\epsilon,\delta}$ (with the corresponding $r_{*}$) gives the same MSE. Hence, the solution $\sigma^{2}=\sigma^{2}_{\epsilon,\delta}$ and $r_{*}=0$ suffices to reach the MMSE when $c=n-1$.
@@
For any $10\leq c<n-$ 1,
$\lim_{\sigma^{2}\to\infty}\min_{r}\min_{\boldsymbol{\alpha}_{\mathcal{U}_{all}}}\mathrm{MSE}(n,\mathcal{U}_{all},\sigma^{2},r,\boldsymbol{\alpha}_{\mathcal{U}_{all}})$
=$\qquad=\begin{cases}O\left(\frac{d\sigma^{2}_{\epsilon,\delta}}{n(n-c)}\right),&\text{if }n(n-c)\gg d\\ O(1),&otherwise.\end{cases}$ (149)
Proof:  Here we prove the claim for the most general case of $c>1$. The proof of other two cases $c=0$ and $c=1$ follow similar steps. Recall that for the case where $|\mathcal{U}|=n$,
$\min_{r}\min_{\boldsymbol{\alpha}_{\mathcal{U}_{all}}}\mathrm{MSE}(n,\mathcal{U}_{all},\sigma^{2},r,\boldsymbol{\alpha}_{\mathcal{U}_{all}})=\left(1+\frac{n/d}{\sigma^{2}+r_{*}(n-1)}\right)^{-1}$ (150)
@@
$r_{*}=\frac{-(n-2)(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})-\sigma^{2}c}{2(n-1)(c-1)}$
$\quad+\frac{\sqrt{((n-2)(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})-\sigma^{2}c)^{2}+4(n-c-1)\sigma^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})}}{2(n-1)(c-1)}$ (151)
@@
×$\quad\times\left(1+\frac{1}{2}\frac{\frac{c^{2}\sigma^{2}\sigma^{2}_{\epsilon,\delta}}{\sigma^{2}-\sigma^{2}_{\epsilon,\delta}}-\sigma^{2}_{\epsilon,\delta}(n-2)^{2}}{\sigma^{2}(n-c)^{2}}\right.$
$\left.\quad-\frac{1}{4}\frac{1}{2!}\left(\frac{\frac{c^{2}\sigma^{2}\sigma^{2}_{\epsilon,\delta}}{\sigma^{2}-\sigma^{2}_{\epsilon,\delta}}-\sigma^{2}_{\epsilon,\delta}(n-2)^{2}}{\sigma^{2}(n-c)^{2}}\right)^{2}+\dotsc\right)$ (153)
@@
$\left.+\sigma^{2}(n\!-\!c)\!\!\left(\!\!1\!-\!\frac{1}{2}\frac{\sigma^{2}_{\epsilon,\delta}}{\sigma^{2}}+\frac{1}{2}\frac{\frac{c^{2}\sigma^{2}\sigma^{2}_{\epsilon,\delta}}{\sigma^{2}-\sigma^{2}_{\epsilon,\delta}}\!-\!\sigma^{2}_{\epsilon,\delta}(n\!-\!2)^{2}}{\sigma^{2}(n-c)^{2}}\!+\!O(\frac{1}{\sigma^{4}})\!\!\right)\!\!\right)$ (154)
$=\frac{1}{2(c-1)}\left(\sigma^{2}_{\epsilon,\delta}(n-2)-\frac{1}{2}\sigma^{2}_{\epsilon,\delta}(n-c)\right.$
$\quad\left.+\frac{c^{2}\sigma^{2}\sigma^{2}_{\epsilon,\delta}}{2(n-c)(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})}-\frac{\sigma^{2}_{\epsilon,\delta}(n-2)^{2}}{2(n-c)}+O(\frac{1}{\sigma^{2}})\right)$ (155)
$=\frac{1}{2(c-1)}\left(\frac{-\sigma^{2}_{\epsilon,\delta}(c-2)^{2}}{2(n-c)}+\frac{c^{2}\sigma^{2}\sigma^{2}_{\epsilon,\delta}}{2(n-c)(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})}\right.$
$\left.\quad+O\left(\frac{1}{\sigma^{2}}\right)\right)$ (156)
@@
$\lim_{\sigma^{2}\to\infty}\min_{r}\min_{\boldsymbol{\alpha}_{\mathcal{U}_{all}}}\mathrm{MSE}(n,\mathcal{U}_{all},\sigma^{2},r,\boldsymbol{\alpha}_{\mathcal{U}_{all}})$
$=\lim_{\sigma^{2}\to\infty}\left(1+\frac{n/d}{\sigma^{2}+r_{*}(n-1)}\right)^{-1}$ (157)
$=\lim_{\sigma^{2}\to\infty}\!\!\!\left(\!\!1\!+\!\frac{n/d}{\frac{1}{2(c-1)}\!\!\!\left(\!\frac{-\sigma^{2}_{\epsilon,\delta}(c-2)^{2}}{2(n-c)}\!+\!\frac{c^{2}\sigma^{2}\sigma^{2}_{\epsilon,\delta}}{2(n-c)(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})}\!+\!O\left(\frac{1}{\sigma^{2}}\right)\right)}\!\!\right)^{\!-1}$ (158)
$=\lim_{\sigma^{2}\to\infty}\left(1+\frac{n(n-c)}{d\sigma^{2}_{\epsilon,\delta}}\right)^{-1}$ (159)
=$=\begin{cases}O\left(\frac{d\sigma^{2}_{\epsilon,\delta}}{n(n-c)}\right),&n(n-c)>>d\\ O(1),&otherwise.\end{cases}$ (160)
@@
Next, we derive the optimum noise parameters for the case with dropouts, i.e., any $\mathcal{U}_{col}$ and $\mathcal{U}$ with $|\mathcal{U}|<n$. We first note that the DME process is the same as what is considered in the case with no dropouts with a reduced number of users. As the noise parameters of the privacy mechanism ($\sigma^{2}$ and $r$) are defined prior to the aggregation stage at the server, they are optimized for the worst case with the largest number of dropouts, resulting in an increase in the MSE compared to the case with no dropouts. However, the decoder can be optimized for the surviving number of users at the time of aggregation, as aggregation is performed after obtaining all responses from the remaining users. Therefore, the proof of the optimum noise parameters and the decoder consists of two steps, namely, 1) determining the optimum decoder for any given number of surviving users, 2) calculating the optimum noise parameters of the privacy mechanism, considering the worst case dropouts and colluding users.
@@
Step 1 is direct from Proposition 1. Proposition 1 provides the optimum decoder for any fixed $\sigma^{2}$ and $r=\rho\sigma^{2}$ with any set of responding users $\mathcal{U}\subseteq\mathcal{U}_{all}$, $|\mathcal{U}|\geq t$. Even though all users may not respond when allowing for dropouts, the same conditions in Lemma 1 must be satisfied by the noise parameters to satisfy the strongest form of $(\epsilon,\delta)$-DP, considering the maximum information leakage that occurs when all $n$ users respond. Therefore, the problem to be solved is given by,
@@
$\min_{r,\sigma^{2}}\max_{\begin{subarray}{c}\mathcal{U}\subset\mathcal{U}_{all}\\ t\leq|\mathcal{U}|<n\end{subarray}}\min_{\boldsymbol{\alpha}_{\mathcal{U}}}\mathrm{MSE}(n,\mathcal{U},\sigma^{2},r,\boldsymbol{\alpha}_{\mathcal{U}})$
@@
$=\min_{\sigma^{2}\geq\sigma^{2}_{\epsilon,\delta}}\!\!\min_{r}\max_{\begin{subarray}{c}\mathcal{U}\subset\mathcal{U}_{all}\\ t\leq|\mathcal{U}|<n\end{subarray}}\left(1+\frac{|\mathcal{U}|/d}{\sigma^{2}+r(|\mathcal{U}|-1)}\right)^{-1}$ (161)
$=\min_{\sigma^{2}\geq\sigma^{2}_{\epsilon,\delta}}\min_{r}\left(1+\frac{t/d}{\sigma^{2}+r(t-1)}\right)^{-1},$ (162)
as $\left(1+\frac{|\mathcal{U}|/d}{\sigma^{2}+r(|\mathcal{U}|-1)}\right)^{-1}$ is decreasing in $|\mathcal{U}|$ for any fixed $\sigma^{2}$ and $r$ that result in a valid covariance matrix, i.e., satisfies $\sigma^{2}+r(n-1)>0$. For any fixed $\sigma^{2}\geq\sigma^{2}_{\epsilon,\delta}$, $\left(1+\frac{t/d}{\sigma^{2}+r(t-1)}\right)^{-1}$ increases with $r$, and the optimal $r$ is given by the $r_{*}$ in (124). The next step is to solve,
$\sigma^{2}_{opt}$ $=\arg\min_{\sigma^{2}\geq\sigma^{2}_{\epsilon,\delta}}\min_{r}\max_{\begin{subarray}{c}\mathcal{U}\subset\mathcal{U}_{all}\\ t\leq|\mathcal{U}|<n\end{subarray}}\min_{\boldsymbol{\alpha}_{\mathcal{U}}}\mathrm{MSE}(n,\mathcal{U},\sigma^{2},r,\boldsymbol{\alpha}_{\mathcal{U}})$ (163)
$=\arg\min_{\sigma^{2}\geq\sigma^{2}_{\epsilon,\delta}}\left(1+\frac{t/d}{\sigma^{2}+r_{*}(t-1)}\right)^{-1}.$ (164)
@@
$\sigma^{2}_{opt}$ $=\frac{\sigma^{2}_{\epsilon,\delta}(n^{2}-2n-cn+2)}{(n-c)^{2}}$
$+\frac{\sigma^{2}_{\epsilon,\delta}(n-c-1)(n+c-2nc+t(n+c-2))}{(n-c)^{2}\sqrt{(t-c)(n-t)(n-c-1)}}.$ (165)
@@
$G(\sigma^{2})$ $=\min_{r}\max_{\begin{subarray}{c}\mathcal{U}\subset\mathcal{U}_{all}\\ t\leq|\mathcal{U}|<n\end{subarray}}\min_{\boldsymbol{\alpha}_{\mathcal{U}}}\mathrm{MSE}(n,\mathcal{U},\sigma^{2},r,\boldsymbol{\alpha}_{\mathcal{U}})$ (166)
$=\left(1+\frac{t/d}{\sigma^{2}+r_{*}(t-1)}\right)^{-1}$ (167)
@@
$\sigma^{2}_{cr}$ $=\frac{\sigma^{2}_{\epsilon,\delta}(n^{2}-2n-cn+2)}{(n-c)^{2}}$
$\quad+\frac{\sigma^{2}_{\epsilon,\delta}\lambda\sqrt{(n-c-1)(n-t)(t-c)}}{(n-c)^{2}(n-t)(t-c)}$ (168)
@@
Proof: $G(\sigma^{2})=\left(1+\frac{t/d}{\sigma^{2}+r_{*}(t-1)}\right)^{-1}$ for any $\sigma^{2}\geq\sigma^{2}_{\epsilon,\delta}$ is a smooth and continuous function in $\sigma^{2}$ as $\sigma^{2}+r_{*}(t-1)$ is smooth and $\sigma^{2}+r_{*}(t-1)>0$ (proof in Section C, Claim 4). Consider the derivative of $G(\sigma^{2})$ with respect to $\sigma^{2}$ denoted by $G^{\prime}(\sigma^{2})$ to find its critical points.
$G^{\prime}(\sigma^{2})\!=\!-\!\left(\!\frac{-\frac{t}{d}\left(1\!+\!(t\!-\!1)\frac{dr_{*}}{d\sigma^{2}}\!\right)}{(\sigma^{2}+r_{*}(t-1))^{2}}\right)\!\!\left(\!1\!+\!\frac{t/d}{\sigma^{2}+r_{*}(t\!-\!1)}\right)^{-2}\!\!\!=0$ (169)
@@
$G^{\prime}(\sigma^{2})$
$=K(t,c,\sigma^{2},r_{*})\left(1+(t-1)\frac{dr_{*}}{d\sigma^{2}}\right)$ (171)
$=K(t,c,\sigma^{2},r_{*})$
×$\times\!\!\left(\!\!1\!+\!\frac{t-1}{2(n-1)(c-1)}\!\left(\!-(n+c-2)+\frac{1}{2}\frac{1}{\sqrt{A(y)}}\frac{dA(y)}{dy}\!\right)\!\right)$ (172)
where $K(t,c,\sigma^{2},r_{*})=\frac{t/d}{(\sigma^{2}+r_{*}(t-1))^{2}}\left(1+\frac{t/d}{\sigma^{2}+r_{*}(t-1)}\right)^{-2}>0$ and
@@
$G^{\prime}(\sigma^{2}_{cr}+\tilde{\delta})$
$=K(t,c,\sigma^{2}_{cr}+\tilde{\delta},r_{*})$
×$\times\left(1+\frac{t-1}{2(n-1)(c-1)}\left(-(n-c-2)+h(\frac{\sigma^{2}_{cr}+\tilde{\delta}}{\sigma^{2}_{\epsilon,\delta}})\right)\right)$ (181)
$=K(t,c,\sigma^{2}_{cr}+\tilde{\delta},r_{*})$
×$\times\left(1+\frac{t-1}{2(n-1)(c-1)}\left(-(n-c-2)+h(y_{cr})+\delta\right)\right)$ (182)
@@
$G^{\prime}(\sigma^{2}_{cr}-\tilde{\delta})$
$=K(t,c,\sigma^{2}_{cr}-\tilde{\delta},r_{*})$
×$\times\left(1+\frac{t-1}{2(n-1)(c-1)}\left(-(n-c-2)+h(\frac{\sigma^{2}_{cr}-\tilde{\delta}}{\sigma^{2}_{\epsilon,\delta}})\right)\right)$ (184)
$=K(t,c,\sigma^{2}_{cr}-\tilde{\delta},r_{*})$
×$\times\left(1+\frac{t-1}{2(n-1)(c-1)}\left(-(n-c-2)+h(y_{cr})-\delta\right)\right)$ (185)
@@
$=K(t,c,\sigma^{2},r_{*})$
$\quad\left(1-\frac{t-1}{n-1}\left(1+\frac{n-2}{(y(n-1)-(n-2))^{2}}\right)\right)$ (189)